# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ 4 Region Of Interest Continuous Ranked Probability Score
where the terms respectively account for CRPS inside the RoI, CRPS outside of the RoI, and constraint satisfaction. We note that the last term is inspired by the threshold-weighted CRPS of Gneiting & Ranjan (2011) and that it vanishes when all constraints are satisfied. The $\alpha$ term is a task-dependent scaling factor that is used to ensure that score values for tasks with numerical data of various scales can be aggregated; its calculation is described in Sec. E.1. Finally, $\beta$ is a scaling factor that controls the impact of constraint violation on the score; we use $\beta=10$ in our experiments. For additional details and discussion on the RCRPR properness, we refer the reader to Appendix E.
@@ 5 Experiments and Results
In this section, we define our evaluation protocol (Sec. 5.1) and outline the baseline models evaluated on CiK (Sec. 5.2). We then present results on the benchmark (Sec. 5.3), along with an analysis of factors affecting model performance. Finally, we look at areas for improvement by analyzing forecasting errors (Sec. 5.4) and inference cost (Sec. 5.4).
@@ 5.1 Evaluation Protocol
Each task in CiK has many unique specifications, i.e. instances arising from the various time series and windows in the associated numerical data, as well as minor variations in natural language context. In order to make the evaluation reproducible and affordable, we deterministically select 5 instances of each task for evaluation. For each instance, we generate 25 independent samples from each model for evaluation. Since many of the tasks in the benchmark share similarities due to having been created from the same data sources or using variants of the same context, we identify these clusters of similar tasks, and design a weighting scheme such that each cluster has equal total weight in our aggregate score (see Sec. A.2 for more details). Finally, to prevent the aggregate scores from being dominated by rare instances where some models give forecasts which are orders of magnitudes away from the ground truth, we introduce an upper bound of 5 to the RCPRS value for each instance, which intuitively represents the value a forecast would get if the distance between the forecast and the ground-truth was 5 times bigger than the range of the ground-truth of the instance.
5.2 Baselines
@@
We evaluate a wide variety of models ranging from methods based on language models to state-of-the-art numerical time series foundation models and classical statistical forecasting methods. Since CiK is meant to be an evaluation benchmark and hence does not have a corresponding training set, we only directly evaluate models that support zero-shot inference (such as LLMs and time series foundation models), and those which can be fit directly to the few historical data points of each task instance evaluated, such as traditional statistical models. We outline these methods below and refer the reader to Appendix D for additional details.
@@
LLM-based Forecasters:  We consider two prompt-based approaches: LLM-processes (LLMP; Requeima et al. (2024)) and a simple approach which we propose, called “Direct Prompt”, where we instruct the model to directly output a forecast as a structured output, rather than prompting it multiple times as in (Requeima et al., 2024) (described in detail in Sec. D.1). For each of these, we evaluate a variety of LLMs with diverse architectures and sizes, such as GPT-4o (Achiam et al., 2023), Mixtral-8x7B (Jiang et al., 2024)), Llama-3-8b (Dubey et al., 2024), and Llama-3.1-405b (Dubey et al., 2024).  [^2] For LLMP, we do not consider Llama-3.1-405b and GPT models as LLMP requires loading model weights into memory, which is infeasible due to resource limitations and confidentiality, respectively. Next, we evaluate multimodal forecasting models, UniTime (Liu et al., 2024b) and Time-LLM (ETTh1) Jin et al. (2024) each trained according to their respective authors’ guidelines (detailed in Sec. D.3). For all of these approaches, inference is performed zero-shot on the benchmark and we compare their performance with and without the natural language context.
@@
Quantitative Forecasting Models:  To contrast the performance of LLM-based forecasters, we also evaluate a number of models that are only capable of processing numerical data (no natural language). This includes exponential smoothing (Gardner Jr., 1985), ETS (Hyndman et al., 2008), and ARIMA (Box et al., 2015), three simple, but time-tested statistical approaches, as well as four state-of-the-art time series foundation models: Lag-Llama (Rasul et al., 2023), Chronos (Ansari et al., 2024)  [^3] Results reported here are on Chronos-Large and Moirai-Large. Results on all versions are in App. C.2. , Moirai (Woo et al., 2024), and TimeGEN (Garza et al., 2023). We note that exponential smoothing, ETS, and ARIMA are fitted to each task instance’s numerical history, while the foundation models are evaluated zero-shot.
Table 1: Results on the CiK benchmark. Starting from the left, the first column shows the RCRPS averaged over all tasks. The second column shows the rank of each method w.r.t. other baselines, averaged over all tasks. The remaining columns show the average RCRPS stratified by model capabilities (Sec. 3.3). All averages are weighted according to the scheme described in Sec. 5.1 and accompanied by standard errors. Lower is better and the best averages are in bold. An asterisk (*) denotes models that do not use natural language context.
Average RCRPS Average Rank Instruction Following Retrieval Reasoning
Direct Prompt (ours)
     Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 4.469 $\pm$ 0.192 0.140 $\pm$ 0.013 0.109 $\pm$ 0.002 0.191 $\pm$ 0.006 0.133 $\pm$ 0.001 0.167 $\pm$ 0.008 0.316 $\pm$ 0.028 0.376 $\pm$ 0.039
     Llama-3-70B-Inst 0.518 $\pm$ 0.030 10.406 $\pm$ 0.190 0.504 $\pm$ 0.038 0.371 $\pm$ 0.071 0.523 $\pm$ 0.048 0.461 $\pm$ 0.048 0.694 $\pm$ 0.117 0.573 $\pm$ 0.044 0.643 $\pm$ 0.049
     Llama-3-8B-Inst 1.647 $\pm$ 0.069 15.130 $\pm$ 0.171 1.604 $\pm$ 0.131 0.199 $\pm$ 0.010 1.568 $\pm$ 0.067 2.133 $\pm$ 0.082 1.555 $\pm$ 0.008 1.589 $\pm$ 0.177 1.840 $\pm$ 0.238
     Mixtral-8x7B-Inst 1.061 $\pm$ 0.058 13.381 $\pm$ 0.230 0.857 $\pm$ 0.077 0.296 $\pm$ 0.049 1.077 $\pm$ 0.078 1.352 $\pm$ 0.117 1.145 $\pm$ 0.144 1.000 $\pm$ 0.086 1.096 $\pm$ 0.106
     GPT-4o 0.276 $\pm$ 0.010 4.368 $\pm$ 0.149 0.180 $\pm$ 0.004 0.087 $\pm$ 0.003 0.519 $\pm$ 0.029 0.113 $\pm$ 0.006 0.447 $\pm$ 0.029 0.590 $\pm$ 0.033 0.769 $\pm$ 0.046
     GPT-4o-mini 0.353 $\pm$ 0.022 8.930 $\pm$ 0.177 0.296 $\pm$ 0.043 0.419 $\pm$ 0.014 0.471 $\pm$ 0.012 0.218 $\pm$ 0.005 1.024 $\pm$ 0.033 0.475 $\pm$ 0.080 0.578 $\pm$ 0.112
LLMP
     Llama-3-70B-Inst 0.550 $\pm$ 0.013 8.038 $\pm$ 0.205 0.645 $\pm$ 0.018 0.284 $\pm$ 0.015 0.392 $\pm$ 0.014 0.519 $\pm$ 0.026 0.312 $\pm$ 0.019 0.453 $\pm$ 0.020 0.495 $\pm$ 0.028
     Llama-3-70B 0.237 $\pm$ 0.006 6.560 $\pm$ 0.254 0.310 $\pm$ 0.011 0.126 $\pm$ 0.009 0.217 $\pm$ 0.007 0.134 $\pm$ 0.003 0.241 $\pm$ 0.019 0.294 $\pm$ 0.008 0.329 $\pm$ 0.010
     Llama-3-8B-Inst 0.484 $\pm$ 0.010 9.457 $\pm$ 0.166 0.345 $\pm$ 0.002 0.138 $\pm$ 0.004 0.910 $\pm$ 0.030 0.242 $\pm$ 0.008 1.278 $\pm$ 0.069 0.617 $\pm$ 0.022 0.787 $\pm$ 0.030
     Llama-3-8B 0.313 $\pm$ 0.023 9.499 $\pm$ 0.323 0.404 $\pm$ 0.043 0.124 $\pm$ 0.003 0.280 $\pm$ 0.026 0.179 $\pm$ 0.014 0.267 $\pm$ 0.015 0.530 $\pm$ 0.084 0.661 $\pm$ 0.117
     Mixtral-8x7B-Inst 0.264 $\pm$ 0.004 8.496 $\pm$ 0.256 0.344 $\pm$ 0.004 0.127 $\pm$ 0.003 0.224 $\pm$ 0.005 0.179 $\pm$ 0.010 0.173 $\pm$ 0.009 0.348 $\pm$ 0.005 0.405 $\pm$ 0.007
     Mixtral-8x7B 0.262 $\pm$ 0.008 8.619 $\pm$ 0.208 0.348 $\pm$ 0.012 0.146 $\pm$ 0.022 0.230 $\pm$ 0.016 0.153 $\pm$ 0.002 0.230 $\pm$ 0.041 0.354 $\pm$ 0.007 0.414 $\pm$ 0.009
Multimodal Models
     UniTime 0.371 $\pm$ 0.002 13.495 $\pm$ 0.091 0.271 $\pm$ 0.003 0.179 $\pm$ 0.001 0.318 $\pm$ 0.001 0.510 $\pm$ 0.003 0.333 $\pm$ 0.001 0.332 $\pm$ 0.001 0.384 $\pm$ 0.001
     Time-LLM (ETTh1) 0.476 $\pm$ 0.001 16.662 $\pm$ 0.075 0.448 $\pm$ 0.002 0.192 $\pm$ 0.000 0.373 $\pm$ 0.000 0.538 $\pm$ 0.003 0.397 $\pm$ 0.001 0.382 $\pm$ 0.001 0.440 $\pm$ 0.001
TS Foundation Models*
     Lag-Llama 0.329 $\pm$ 0.004 13.157 $\pm$ 0.235 0.355 $\pm$ 0.007 0.181 $\pm$ 0.003 0.324 $\pm$ 0.003 0.272 $\pm$ 0.006 0.342 $\pm$ 0.006 0.386 $\pm$ 0.009 0.449 $\pm$ 0.012
     Chronos 0.326 $\pm$ 0.002 11.962 $\pm$ 0.139 0.385 $\pm$ 0.002 0.138 $\pm$ 0.002 0.288 $\pm$ 0.002 0.249 $\pm$ 0.002 0.295 $\pm$ 0.003 0.362 $\pm$ 0.003 0.417 $\pm$ 0.004
     TimeGEN 0.354 $\pm$ 0.000 14.345 $\pm$ 0.090 0.402 $\pm$ 0.000 0.176 $\pm$ 0.000 0.308 $\pm$ 0.000 0.279 $\pm$ 0.000 0.324 $\pm$ 0.000 0.377 $\pm$ 0.000 0.431 $\pm$ 0.000
     Moirai 0.520 $\pm$ 0.006 12.458 $\pm$ 0.266 0.414 $\pm$ 0.004 0.155 $\pm$ 0.004 0.260 $\pm$ 0.003 0.751 $\pm$ 0.015 0.276 $\pm$ 0.008 0.337 $\pm$ 0.007 0.397 $\pm$ 0.010
Statistical Models*
     ARIMA 0.480 $\pm$ 0.006 12.320 $\pm$ 0.180 0.399 $\pm$ 0.006 0.160 $\pm$ 0.002 0.517 $\pm$ 0.012 0.522 $\pm$ 0.013 0.706 $\pm$ 0.026 0.354 $\pm$ 0.007 0.403 $\pm$ 0.010
     ETS 0.522 $\pm$ 0.009 14.310 $\pm$ 0.203 0.407 $\pm$ 0.009 0.228 $\pm$ 0.010 0.682 $\pm$ 0.018 0.571 $\pm$ 0.019 0.855 $\pm$ 0.035 0.453 $\pm$ 0.012 0.479 $\pm$ 0.015
     Exp-Smoothing 0.603 $\pm$ 0.013 14.936 $\pm$ 0.132 0.571 $\pm$ 0.021 0.334 $\pm$ 0.013 0.743 $\pm$ 0.018 0.557 $\pm$ 0.019 0.899 $\pm$ 0.035 0.673 $\pm$ 0.038 0.782 $\pm$ 0.053
5.3 Results on the benchmark
@@
Our main results are shown in Tab. 1. At a high level, we observe that the best-performing baselines combine pretrained LLMs with prompting strategies like Direct Prompt and LLMP, with a bias toward the largest models. In terms of RCRPS, Llama-3.1-405B-Inst (Direct Prompt) significantly outperforms all of its counterparts. As can be seen in Fig. 4, it achieves this only with context. GPT-4o (Direct Prompt) performs worse with respect to RCRPS, but compares favorably in terms of average rank, taking the best average rank by a small margin. This discrepancy is due to strong failures on some of the tasks, which we discuss in Sec. 5.4. Other models like Llama-3-70B (LLMP), Mixtral-8x7B-Inst (LLMP), Mixtral-8x7B (LLMP), and Llama-3-8B (LLMP) are on par with GPT-4o (Direct Prompt) in terms of RCRPS. Interestingly, all of these baselines outperform UniTime and Time-LLM, which also rely on LLMs (GPT-2 & LLaMA-7B). We discuss this gap in Tab. 8. Finally, as emphasized in Fig. 5, we observe that the best-performing LLM baselines significantly outperform purely quantitative models. In what follows, we examine various aspects of these results.
Explaining the performance of
LLM-based approaches
Figure 4: Performance with and without context (lower is better). Full bars show performance with context; striped bars show performance without. All models improve with context, except DP Mixtral-8x7B-Inst, LLMP Llama-3-70B-Inst and Time-LLM. Llama-3.1-405B-Inst improves significantly with context, exhibiting the best aggregate RCRPS.
The stronger performance of LLM baselines could be due to two factors: (i) properly leveraging the natural language context and (ii) being more proficient at numerical forecasting. We thus attempt to disentangle their contributions. On the one hand, Fig. 4 shows clear evidence that most baselines make use of the context to improve their forecasts. For example, Llama-3.1-405B-Inst (Direct Prompt) improves by $67.1\%$ with context. This is reflected in the quality of the forecasts, where we observe clear improvements especially in regions of interest and improved satisfaction of constraints (see Sec. C.4 for examples). Other models show much slighter improvements and, in three cases, even a degradation in performance. These can be explained either by the context being ignored, or by significant failures in using context, impoverishing overall performance (see Sec. 5.4).
@@
Figure 5: Proportion of tasks for which LLM-based baselines outperform the 7 quantitative forecasting baselines (see Sec. 5.2). A baseline is considered to outperform another on a task if its mean RCPRS is lower on said task. Results are shown for variants that use (left) and do not use (right) the natural language context. A full green bar would indicate that the baseline is better on all tasks, whereas a full red bar would indicate that it is worse everywhere. Averages are weighted according to Sec. 5.1.
@@
On the other hand, Fig. 5 (right) shows that some LLM baselines are surprisingly good forecasters when compared to quantitative forecasting models in a no-context setting. For instance, multiple Llama-3-based models used with the LLMP strategy outperform at least $5$ of the quantitative baselines on the majority of tasks. This is further substantiated by results in Sec. C.2. In contrast, other baselines, including the best models Llama-3.1-405B-Inst (Direct Prompt) and GPT-4o (Direct Prompt), show much weaker numerical forecasting abilities without context, suggesting that their performance is mostly due to leveraging the context.
Comparing the LLMP and Direct Prompting Strategies
@@
Clear patterns emerge when comparing these strategies. First, as shown in Fig. 5 (right), LLMP baselines exhibit stronger numerical forecasting performance without context than Direct Prompt baselines. This advantage likely stems from LLMP’s closer alignment with the forecasting task: LLMP simply prompts the LLM to autoregressively predict the next value in the time series – a task well suited for non-instruction tuned LLMs. This contrasts with Direct Prompting which requires output forecasts to be structured, complicating the overall task.
@@
This line of reasoning leads us to our second observation; as reflected in Tab. 1 and Fig. 5, instruction tuning appears to degrade LLMP performance, with Llama-3 models showing a twofold decrease in performance after tuning—a behavior previously observed by Gruver et al. (2024). Interestingly, instruction tuning does not degrade Mixtral-8x7B performance. Finally, while instruction tuning generally harms LLMP, it is essential for models used with the Direct Prompt strategy. Again, Direct Prompt requires forecasts to be produced in a specific structure, a skill that base models typically hone during post-training (see Sec. D.1.1 for details).
No Baseline Excels Across All Capabilities
@@
Based on the results in Tab. 1, it is evident that some models possess the necessary capabilities to effectively utilize the contextual information provided. However, we observe that no single model is the best across all capabilities. Llama-3.1-405B-Inst (Direct Prompt), our overall top-performing baseline, outperforms its counterparts in only 4 out of 7 capabilities. This finding indicates that the benchmark remains unsolved, leaving significant room for advancements from the research community.

### v2

@@ 4 Region of Interest CRPS
where the terms respectively account for the CRPS inside the RoI, the CRPS outside of the RoI, and the constraint violation penalty. The last term, which is inspired by the threshold-weighted CRPS of Gneiting & Ranjan (2011), vanishes when all constraints are satisfied. The $\alpha$ term is a task-dependent normalization factor to make the RCRPS scale-independent, which enables fair RCRPS aggregation across tasks; its calculation is described in Sec. E.1. Finally, $\beta$ is a scaling factor that controls the impact of constraint violation on the score; we use $\beta=10$ in our experiments. We refer the reader to Appendix E for additional details.
@@ 5 Experiments and Results
In this section, we define our evaluation protocol (Sec. 5.1) and outline the models that we evaluate on CiK (Sec. 5.2). We then present results on the benchmark (Sec. 5.3), along with an analysis of factors affecting model performance. Finally, we look at areas for improvement by analyzing forecasting errors (Sec. 5.4) and inference cost (Sec. 5.5).
@@ 5.1 Evaluation Protocol
Each task in CiK has many unique instances arising from the selection of time series and windows in the associated numerical data, as well as minor variations in natural language context. We deterministically sample five instances of each task in order to make the evaluation reproducible and affordable. For every instance, we generate 25 independent forecasts per model for evaluation. Many of the tasks in the benchmark share similarities due to sharing the same data sources or using variants of the same context. Therefore, we identify clusters of similar tasks and design a weighting scheme such that each cluster has equal total weight in our aggregate score (see Sec. A.5 for more details).
5.2 Methods
@@
We evaluate a wide variety of models including methods based on LLMs, state-of-the-art numerical time series foundation models (TSFMs) and classical statistical forecasting methods. Since CiK is meant to be an evaluation benchmark, it does not have a corresponding training set. We therefore only evaluate models that can produce forecasts directly based on the history of a series. This includes LLMs and TSFMs that support zero-shot inference, and traditional statistical models that can be fit directly to the history of a given instance. We outline these methods below and refer the reader to Appendix D for additional details.
@@
LLM-based Forecasters:  We consider two prompt-based approaches. We propose “ direct prompt ”, a simple approach where we instruct the model to directly output a forecast as a structured output for all of the required timestamps (see Sec. D.1 for more details). We also evaluate LLM Processes (llmp; Requeima et al. (2024)), a method which autoregressively prompts the LLM multiple times to output a forecast. For each of these, we evaluate a variety of LLMs with diverse architectures and sizes, such as GPT-4o , GPT-4o-mini (Achiam et al., 2023), Qwen-2.5-{0.5B, 1.5B, 7B}  (Yang et al., 2024), Mixtral-8x7B (Jiang et al., 2024), Llama-3-{8B, 70B}  (Dubey et al., 2024), Llama-3.1-405B (Dubey et al., 2024). [^1] For llmp, we do not consider Llama-3.1-405B and GPT models as llmp requires loading model weights into memory, which is infeasible due to resource limitations and confidentiality of the respective models (see Sec. D.2.2 for more details). We also evaluate multimodal forecasting models, UniTime  (Liu et al., 2024c) and Time-LLM (ETTh1)  (Jin et al., 2024) each trained according to their respective authors’ guidelines (details in Sec. D.3). For all of these approaches, inference is performed zero-shot on the benchmark and we compare their performance with and without the natural language context.
@@
Quantitative Forecasting Models:  We also evaluate a number of models that process only numerical data, but cannot integrate natural language. We evaluate Exponential Smoothing  (Gardner Jr. (1985)), ETS  (Hyndman et al., 2008), and ARIMA  (Box et al., 2015), three simple, but time-tested statistical approaches. We also evaluate four state-of-the-art TSFMs: Lag-Llama  (Rasul et al., 2023), Chronos  (Ansari et al., 2024), Moirai  (Woo et al., 2024), and TimeGEN  (Garza et al., 2023). We fit Exponential Smoothing , ETS and ARIMA to the individual numerical history of each task instance, which is the same input that the TSFMs process to produce forecasts zero-shot.
5.3 Results on CiK
@@
Tab. 1 shows our main results. At a high level, we observe that the best-performing methods combine pretrained LLMs with prompting strategies like direct prompt and llmp, with a bias towards the largest models. In terms of RCRPS, Llama-3.1-405B-Inst ( direct prompt ) significantly outperforms all of its counterparts. GPT-4o ( direct prompt ) performs worse with respect to RCRPS, but compares favorably in terms of average rank. This discrepancy is due to strong failures on some of the tasks, which we discuss in Sec. 5.4. Other models like Llama-3-70B (llmp), Mixtral-8x7B-Inst (llmp) and Mixtral-8x7B (llmp) are on par with GPT-4o ( direct prompt ), Llama-3-70B-Inst ( direct prompt ) and Qwen-2.5-7B-Inst ( direct prompt ) in terms of RCRPS. Interestingly, all of these methods outperform UniTime and Time-LLM , which also rely on LLMs (GPT-2 & LLaMA-7B). We discuss this gap in Tab. 10. Finally, as emphasized in Figure 5, we observe that the best-performing LLM methods significantly outperform purely quantitative models. In what follows, we examine various aspects of these results (and refer to Appendix C for additional results).
@@
Figure 5: Proportion of tasks for which LLM-based methods outperform the 7 quantitative forecasting methods (see Sec. 5.2). A method is considered to outperform another on a task if its average RCPRS is lower on said task. Results are shown for variants that use (left) and do not use (right) the natural language context. A full green bar would indicate that the method is better on all tasks, whereas a full red bar would indicate that it is worse everywhere. Tasks are weighted according to Sec. 5.1.
Explaining the performance of LLM-based approaches:  The strong performance of LLM-based methods could be due to two factors: (i) properly leveraging the natural language context and (ii) being more proficient at numerical forecasting. We thus aim to disentangle their contributions.
On the one hand, Figure 6 shows clear evidence that methods with access to the context improve their forecasts. For example, Llama-3.1-405B-Inst ( direct prompt ) improves by $67.1\%$ with context. We find these differences statistically significant across many of the models (see Sec. C.6). This is reflected in the quality of the example forecasts in Sec. C.11, where we observe clear improvements in regions of interest, as well as improved constraint satisfaction. Other models show lesser improvements and, in some cases, even a degradation in performance. Our analysis in Sec. 5.4 shows that this can be explained either by the context being ignored, or by significant failures in using context, worsening overall performance.
@@
On the other hand, Figure 5 (right) shows that LLM-based forecasters, when evaluated without context, no longer dominate the quantitative forecasting models. However, some LLM-based forecasters remain surprisingly competitive. For instance, multiple Llama-3 (llmp) models outperform at least $5$ of the quantitative models on the majority of tasks. The extended results in Sec. C.1 further substantiate this. In contrast, models such as Llama-3.1-405B-Inst ( direct prompt ) and GPT-4o ( direct prompt ) show significantly weaker forecasting performance without context. This suggests that such models are especially preferable in cases where context is available. This is also reflected in their aggregate scores without context (in App. Tab. 2).
@@
Comparing llmp and direct prompt :  Clear patterns emerge when comparing these strategies. Figure 5 (right) shows that without context, llmp models exhibit stronger numerical forecasting performance than direct prompt models. This advantage likely stems from llmp ’s closer alignment with the forecasting task: llmp simply prompts the LLM to autoregressively predict the next value in the time series, a task well suited to base models with no instruction tuning. In contrast, direct prompt requires output forecasts to be structured, which relatively complicates the task.
@@
This line of reasoning leads us to another observation on the impact of instruction tuning; as reflected in Tab. 1 and Figure 5, instruction tuning appears to generally degrade llmp performance. Llama-3 models showing a twofold decrease in performance after tuning—a behavior previously observed by Gruver et al. (2024). Interestingly, instruction tuning does not degrade the performance of Mixtral-8x7B . Finally, whereas llmp mostly suffers from instruction tuning, direct prompt requires forecasts to be produced in a specific structure, a skill that instruction-tuned models might be better at (see Sec. D.1.1 for details).
@@
No Method Excels Across All Context Types:  Tab. 1 shows that some methods can effectively produce better forecasts with the provided contextual information. However, no single method is the best across context types. Llama-3.1-405B-Inst ( direct prompt ), the top-performing method on average, outperforms its counterparts on only 3 out of 5 context types. This finding indicates that the benchmark remains unsolved, leaving significant room for advancements from the research community.
Figure 6: RCRPS with and without context (log scale, lower is better). Full bars show performance with context; striped bars show performance without. In general, larger models outperform smaller models and benefit much more from context. direct prompt models all improve with context, other than Qwen-2.5-0.5B-Instruct . For llmp, larger models benefit from context, but smaller models fail to do so and perform worse in general.